# Nautilus: Locality-aware Autoencoder for Scalable Mesh Generation

### v4

@@ Abstract.
Triangle meshes are fundamental to 3D applications, enabling efficient modification and rasterization while maintaining compatibility with standard rendering pipelines. However, current automatic mesh generation methods typically rely on intermediate representations that lack the continuous surface quality inherent to meshes. Converting these representations into meshes produces dense, suboptimal outputs. Although recent autoregressive approaches demonstrate promise in directly modeling mesh vertices and faces, they are constrained by the limitation in face count, scalability, and structural fidelity. To address these challenges, we propose Nautilus, a locality-aware autoencoder for artist-like mesh generation that leverages the local properties of manifold meshes to achieve structural fidelity and efficient representation. Our approach introduces a novel tokenization algorithm that preserves face proximity relationships and compresses sequence length through locally shared vertices and edges, enabling the generation of meshes with an unprecedented scale of up to 5,000 faces. Furthermore, we develop a Dual-stream Point Conditioner that provides multi-scale geometric guidance, ensuring global consistency and local structural fidelity by capturing fine-grained geometric features. Extensive experiments demonstrate that Nautilus significantly outperforms state-of-the-art methods in both fidelity and scalability. The project page is at https://nautilusmeshgen.github.io.
copyright: noneccs: Computing methodologies Mesh geometry modelsccs: Computing methodologies Mesh modelsccs: Computing methodologies Shape inference
@@
Figure 1. Capability highlight of our Nautilus: Given a point cloud (e.g., the bird, aircraft, penguin, Squog, shield, and bowl shown in the figure) or a single image (e.g., the gorilla, ax, man, planet, sycee, and bear shown in the figure) as input, Nautilus enables the direct generation of aesthetic, artist-like mesh assets conditioned on the input. By explicitly modeling and leveraging locality, it delivers superior generation quality and unprecedentedly detailed topologies.
**footnotetext: These authors contributed equally to this work.
1. Introduction
@@
Triangle meshes, composed of interconnected triangles that approximate arbitrary topology in space, are fundamental to 3D applications (Dai and Nießner, 2019; Siddiqui et al., 2024; Hong et al., 2024). Their structure enables efficient modification and rasterization, ensuring compatibility with standard rendering pipelines. Currently, creating high-quality meshes primarily relies on skilled 3D artists with tools such as Blender, making the process labor-intensive and difficult to scale. Therefore, automating high-quality mesh generation is essential to meet the demands of modern 3D applications.
@@
However, current automatic approaches mainly adopt intermediate representations such as Voxels (Ren et al., 2024; Wu et al., 2016; Brock et al., 2016), Point cloud (Luo and Hu, 2021; Nichol et al., 2022; Zhou et al., 2021), NeRF (Poole et al., 2022; Wang et al., 2024; Yi et al., 2024b; Hong et al., 2023), and 3DGS (Tang et al., 2025; Shen et al., 2024; Yi et al., 2024a; Zhang et al., 2024a; Wang et al., 2025), which inherently lack the continuous surface of the meshes. Converting these representations into meshes through marching cubes (Lorensen and Cline, 1998) or Poisson reconstruction (Kazhdan et al., 2006) often yields dense, suboptimal meshes, lacking the conciseness and aesthetic of artist-created models. Even with post-processing such as remeshing (Attene and Falcidieno, 2006), the generated meshes remain substantially inferior to artist-created assets, as highlighted by (Chen et al., 2024b) and (Siddiqui et al., 2024).
@@
Very recently, a line of works has showcased the potential of directly modeling mesh structures (Siddiqui et al., 2024; Chen et al., 2024b, a) in an auto-regressive manner. Pioneered by MeshGPT (Siddiqui et al., 2024), this paradigm tokenizes mesh faces into sequences and employs a GPT-style decoder to autoregressively generate the sequences, enabling direct capture of vertex-face relationships. MeshAnything (Chen et al., 2024b) further enhanced this framework by incorporating point cloud conditioning for global geometric guidance. Through explicit modeling of mesh vertices and faces, these methods generate more concise and structurally refined meshes.
@@
Despite these advances, current direct mesh generation still encounters several fundamental challenges: (1) Existing methods often struggle to maintain local structure fidelity (Weng et al., 2024; Chen et al., 2024b, c), as shown in Fig. 2(a), exhibiting manifold defects such as surface holes, overlapping faces, and missing components, especially when dealing with complex topologies. (2) Moreover, existing methods are typically limited to modeling meshes with fewer than 1,600 faces, which significantly constrains their ability to capture the intricate topological details required for industrial applications, as illustrated in Fig. 2(b).
@@
Drawing from 3D artists’ creation with precise structure and seamless transitions, we conjecture that the crux arises from inadequate modeling of manifold mesh locality—where adjacent faces share common edges and each face cluster converges at a central vertex. This locality offers two key inspirations to overcome the aforementioned challenges: (1) Local structure preservation. The locality ensures that the geometry of each face is directly constrained by its immediate neighbors. Thus, prioritizing local dependencies on neighbors and their topology could be crucial for faces to form precise interconnection, achieving manifoldness and structure fidelity rather than merely approximating the shape contour. (2) Efficient mesh representation. Explicitly modeling the locally shared edges and vertices could significantly reduce redundancy in representation and compress the sequence, allowing the modeling of refined meshes with higher face counts and more complex topology.
@@
Figure 2. Fundamental challenges in current methods. (a) Being restricted to generating meshes with fewer faces leads to the loss of fine topological details compared to higher face count meshes. (b) Struggling to preserve local structure fidelity causes manifold defects like surface holes and overlapping faces, in contrast to well-structured assets.
@@
Figure 3. The overall pipeline of Nautilus. Given a point cloud or image as input, its transformer decoder autoregressively generates mesh sequence tokenized by the Nautilus-style algorithm. Through detokenization, the output tokens are converted to interconnected mesh faces, constructing our generated mesh assets.
2. Related Work
@@
Mesh Generation with Intermediate Representation. Most existing 3D generation methods create meshes through intermediate representations, including neural fields (Poole et al., 2022; Yi et al., 2024b; Wu et al., 2024b), 3DGS (Tang et al., 2023; Chung et al., 2023), and SDF (Park et al., 2019; Xu et al., 2024). Utilizing SDS optimization (Poole et al., 2022; Lin et al., 2023; Wang et al., 2024) with pre-trained 2D models (Rombach et al., 2022; Esser et al., 2024), early approaches achieved high quality but required hours of computation per instance. Following the emergence of large-scale 3D datasets, research shifted toward feed-forward 3D models (Hong et al., 2023; Yi et al., 2024a; Tang et al., 2025). Pioneered by the Large Reconstruction Model (LRM), one line of work (Shen et al., 2024; Zhang et al., 2024a) demonstrated that end-to-end training with differentiable rendering and Transformer-based models (Vaswani, 2017) could generate 3D assets within seconds. Another line of work (Zhang et al., 2024b; Li et al., 2024b; Wu et al., 2024a) employs 3D-native latent diffusion with volumetric representation, encoding 3D assets into compressed latent spaces for generation through diffusion steps. However, these methods require decent post-processing (Lorensen and Cline, 1998; Attene and Falcidieno, 2006) and often produce overly dense or over-smooth meshes.
@@
Direct Mesh Generation. Compared to methods using intermediate representations, direct mesh generation approaches explicitly model mesh topology, producing more concise and structurally coherent meshes without post-conversion steps. Although pioneering works such as PolyGen (Nash et al., 2020) and PolyDiff (Alliegro et al., 2023) demonstrated promise, their applications remained limited to single-category datasets. Recent advances in autoregressive approaches, notably MeshGPT (Siddiqui et al., 2024), introduced mesh tokenization with VQ-VAE (Van Den Oord et al., 2017) compression for direct mesh generation under topology supervision. Subsequent research has advanced this paradigm by exploring various architectural variants (Chen et al., 2024a; Weng et al., 2024) and extending it to conditional generation tasks (Chen et al., 2024b, c), particularly point cloud to mesh generation. However, these methods exhibit manifold defects when handling complex topologies and are limited to 1,600 faces due to the excessive length of tokenized sequence. In contrast, our Nautilus aims to achieve manifold structure fidelity through efficient locality-aware mesh representation while extending the maximum face count.
@@
Figure 4. The illustration of our Nautilus-style Mesh Tokenization algorithm. (a) The shell-structured traversal of mesh faces. (b) Top: The continuous shell traversal mechanism designed to find the center of the next shell. Bottom: The tokenized vertices sequence of the shown example.
3. Nautilus
@@
In this section, we introduce Nautilus, a locality-aware autoencoder for artist-like mesh generation, with the overall pipeline depicted in Fig. 3. Initially, we design a novel Nautilus-style algorithm that tokenizes artist-created meshes into sequences (Sec. 3.1), which serves as the training input of our framework. During generation, Nautilus takes either a point cloud or a single-view image as the condition inputs. For point cloud conditioning, we propose a dual-stream conditioner (Sec. 3.2) that provides both global and local geometric guidance to the autoregressive decoder, which generates the tokenized sequence given the input condition (Sec. 3.3). For image conditioning, we integrate a pre-trained image encoder with the decoder to enable image-conditioned generation (Sec. 3.4). Finally, after detokenization, the token sequences predicted by the decoder are translated into mesh faces, constructing our generated artist-like mesh assets.
3.1. Nautilus-style Mesh Tokenization
@@
The initial step of autoregressive mesh generation involves serializing the triangle mesh into tokens (Chen et al., 2024a; Siddiqui et al., 2024; Chen et al., 2024b). A mesh asset $\mathbf{M}=\{\mathbf{f}^{i}\}_{i=1}^{N}$ consists of $N$ triangle faces $\mathbf{f}^{i}$, each of which is represented by three vertices, $\{\mathbf{v}_{1}^{i},\mathbf{v}_{2}^{i},\mathbf{v}_{3}^{i}\}$. By flattening the discrete 3D coordinates of each vertex, denoted as $\mathbf{v}_{j}^{i}=\{x_{j}^{i},y_{j}^{i},z_{j}^{i}\}$, the mesh asset $\mathbf{M}$ can be serialized into an ordered sequence of coordinates, $S(\mathbf{M})$, with a total length of $9N$:
(1) $S(\mathbf{M})=\{\mathbf{v}_{1}^{1},\mathbf{v}_{2}^{1},\mathbf{v}_{3}^{1},\mathbf{v}_{1}^{2},\mathbf{v}_{2}^{2},\mathbf{v}_{3}^{2},\dots,\mathbf{v}_{1}^{N},\mathbf{v}_{2}^{N},\mathbf{v}_{3}^{N}\}$
@@
Nautilus Shell Representation. In tokenization, we first sort the mesh faces by their coordinates $z$-$y$-$x$ to facilitate efficient serialization. As depicted in Fig. 4(a), we divide the mesh asset into multiple shells for sequential generation. Each shell organizes faces around a central vertex $O$ with an ordered sequence of surrounding vertices $P$, where each face is formed by $O$ and two adjacent $P$ as $\mathbf{f}_{OP_{i}P_{i+1}}$. This organization helps maintain local mesh connectivity by explicitly encoding edge-sharing relationships between adjacent faces. Based on this shell structure, we compactly represent $N$ faces using a sequence of $(N+2)$ vertices, rather than the $3N$ vertices required in Eq. (1):
(2) $S(\mathbf{f}_{OP_{1}P_{2}},\mathbf{f}_{OP_{2}P_{3}},\dots)=\{O,P_{1},P_{2},P_{3},\dots\}$
@@
With this representation, the shell #1 in Fig. 4(a) can be serialized to $\{O_{1},P_{1,1},P_{1,2},\dots,P_{1,10},P_{1,1}\}$, which preserves the adjacency of neighboring vertices in sequence. By explicitly modeling the locally shared vertices and faces, we could add a new face to a shell by extending the sequence with only a single vertex.
@@
Continuous Shell Traversal. After completing the traversal of one shell, we decide the central vertex candidate for the next shell based on two insights: (1) The vertex should be close to the current shell to facilitate local, short-distance dependencies and promote coherent edge connectivity between shells. (2) The vertex should have a relatively high degree and is shared by more faces, which not only improves the compression ratio but also encourages the formation of well-connected local mesh structures. Therefore, we design the following continuous shell traversal strategy. Specifically, after completing one shell, we list all neighbor vertices of its last traversed vertex and select the neighbor with maximum degree as the center vertex of the next shell  [^1] As depicted in Fig. 4(b), the last vertex of the shell #k is located at $P_{k,\text{end}}$, therefore we choose the orange vertex with the highest degree among its neighbors as the next center $O_{k+1}$.. Our strategy enhances the traversal continuity and spatial proximity between shells, promoting locally consistent mesh topology.
@@
Coordinates Compression. Building upon the vertex sharing optimization in our Nautilus representation, we exploit spatial quantization and dimensional reduction techniques to achieve more compact mesh encoding. Specifically, given the 3D spatial resolution $\alpha$, we can uniquely represent the discrete coordinates $(x,y,z)\in\mathbb{R}^{\alpha}$ as $(u,v)$ using a scaling factor $\beta$:
(3) $x*\alpha^{2}+y*\alpha+z=u*\beta+v$
@@
By mapping the $x$-$y$-$z$ coordinates to $u$-$v$ space, we could further compress the sequence length. With a resolution of 128 and a scaling factor of 2048, we construct a codebook of size 1024 for $u$ and 2048 for $v$. Additionally, to identify the start of each shell, we extend the codebook by 1024 to separately encode the $u$ coordinates of center vertices $O$, thereby distinguishing them from surrounding vertices $P$. This design avoids the increased sequence length introduced by the special tokens that previous work (Chen et al., 2024c; Tang et al., 2024) used for sub-sequence separation, producing our final tokenized sequence as:
(4) $S(\mathbf{M})=\{u^{\text{O}}_{1},v^{\text{O}}_{1},u^{\text{P}}_{1,1},v^{\text{P}}_{1,1},u^{\text{P}}_{1,2},\dots,u^{\text{O}}_{2},v^{\text{O}}_{2},\dots\}$
@@
In a nutshell, our Nautilus-style tokenization provides a compact serialization to model complex mesh assets with a higher number of faces. Meanwhile, it maintains short-distance local dependencies by preserving the proximity of spatially neighboring vertices in the tokenized sequence, enhancing the fidelity of local structure. Experiments in Sec. 4 show that these advantages lead to significant improvements in scalability and generation quality. A detailed step-by-step demonstration of the tokenization process is available in Appendix, further illustrating its design and implementation.
3.2. Dual-Stream Point Conditioner

### v5

@@ Abstract
Triangle meshes are fundamental to 3D applications. Current automatic mesh generation methods typically rely on intermediate representations that lack the continuous surface quality inherent to meshes. Converting these representations into meshes produces dense, suboptimal outputs. Although recent autoregressive approaches demonstrate promise in directly modeling mesh vertices and faces, they are constrained by the limitation in face count, scalability, and structural fidelity. To address these challenges, we propose Nautilus, a locality-aware autoencoder for artist-like mesh generation that leverages the local properties of manifold meshes to achieve structural fidelity and efficient representation. Our approach introduces a novel tokenization algorithm that preserves face proximity and compresses sequence length through locally shared vertices and edges, enabling the generation of meshes with an unprecedented scale of up to 5,000 faces. Furthermore, we develop a Dual-stream Point Conditioner that captures fine-grained geometry, ensuring global consistency and local structural fidelity. Our experiments demonstrate that Nautilus significantly outperforms existing methods in generation quality.
@@
Figure 1: Given a point cloud (e.g., the bird, aircraft, penguin, Squog, shield, and bowl shown in the figure) or a single image (e.g., the gorilla, ax, man, planet, sycee, and bear shown in the figure) as input, Nautilus enables the direct generation of mesh assets.
1 Introduction
@@
Triangle meshes, composed of interconnected triangles that approximate arbitrary topology in space, are fundamental to 3D applications [10, 31, 15, 8, 14, 12, 13]. Their structure enables efficient modification and rasterization, ensuring compatibility with standard rendering pipelines. Currently, creating high-quality meshes primarily relies on skilled 3D artists, making the process labor-intensive and difficult to scale. Therefore, automating high-quality mesh generation is essential to meet the demands of modern 3D applications.
@@
However, current automatic approaches mainly adopt intermediate representations such as Voxels [27, 41, 4], Point cloud [22, 24, 51], NeRF [26, 39, 46, 16], and 3DGS [34, 30, 45, 48, 38], which inherently lack the continuous surface of the meshes. Converting these representations into meshes through marching cubes [21] or Poisson reconstruction [17] often yields dense, suboptimal meshes, lacking the conciseness and aesthetic of artist-created models. Even with post-processing such as remeshing [2], the generated meshes remain substantially inferior to artist-created assets, as highlighted by [6] and [31].
@@
Very recently, a line of works has showcased the potential of directly modeling mesh structures [31, 6, 5] in an auto-regressive manner. Pioneered by MeshGPT [31], this paradigm tokenizes mesh faces into sequences and employs a GPT-style decoder to autoregressively generate the sequences, enabling direct capture of vertex-face relationships. MeshAnything [6] further enhanced this framework by incorporating point cloud conditioning for global geometric guidance. Through explicit modeling of mesh vertices and faces, these methods generate more concise and structurally refined meshes.
@@
Despite these advances, current direct mesh generation still encounters several fundamental challenges: (1) Existing methods often struggle to maintain local structure fidelity [40, 6, 7], as shown in Figure 2(a), exhibiting manifold defects such as surface holes, overlapping faces, and missing components, especially when dealing with complex topologies. (2) Moreover, most existing methods are typically limited to modeling meshes with insufficient face number, which significantly constrains their ability to capture the intricate topological details required for industrial applications, as illustrated in Figure 2(b).
@@
Drawing from artists’ creation with precise structure and seamless transitions, we conjecture that the crux arises from inadequate modeling of manifold mesh locality—where adjacent faces share common edges and each face cluster converges at a central vertex. This locality offers two key inspirations to overcome the aforementioned challenges: (1) Local structure preservation. The locality ensures that each face’s geometry is directly constrained by its immediate neighbors. Thus, prioritizing local dependencies on neighbors and their topology could be crucial for faces to form precise interconnection, achieving manifoldness and structure fidelity rather than merely approximating the shape contour. (2) Efficient mesh representation. Explicitly modeling the locally shared edges and vertices could significantly reduce redundancy in representation and compress the sequence, allowing the modeling of refined meshes with higher face counts and more complex topology.
@@
Figure 2: Fundamental challenges in current methods. (a) Being restricted to generating meshes with fewer faces leads to the loss of fine topological details. (b) Struggling to preserve local structure fidelity causes manifold defects like holes and overlapping faces.
@@
Figure 3: The overall pipeline of Nautilus. Given a point cloud or image as input, its transformer decoder autoregressively generates mesh sequence tokenized by the Nautilus-style algorithm. Through detokenization, the output tokens are converted to interconnected mesh faces.
2 Related Work
@@
Mesh Generation with Intermediate Representation. Most existing 3D generation methods create meshes through intermediate representations, including neural fields [26, 46, 43], 3DGS [32, 9], and SDF [25, 44]. Utilizing SDS optimization [26, 20, 39] with pre-trained 2D models [28, 11], early approaches achieved high quality but required hours of computation per instance. Following the emergence of large-scale 3D datasets, research shifted toward feed-forward 3D models [16, 45, 34]. Pioneered by the Large Reconstruction Model (LRM), one line of work [30, 48] demonstrated that end-to-end training with differentiable rendering and Transformer-based models [37] could generate 3D assets within seconds. Another line of work [49, 18, 42] employs 3D latent diffusion with volumetric representation, encoding 3D assets into compressed latent spaces for generation through diffusion steps. However, these methods require decent post-processing [21, 2] and often produce overly dense or over-smooth meshes.
@@
Direct Mesh Generation. Compared to methods using intermediate representations, direct mesh generation approaches explicitly model mesh topology, producing more concise and structurally coherent meshes without post-conversion steps. Pioneering works such as PolyGen [23] and PolyDiff [1] showed promise but remained limited to single-category datasets. Recent advances in autoregressive approaches, notably MeshGPT [31], introduced mesh tokenization with VQ-VAE [36] compression for direct mesh generation under topology supervision. Subsequent research has advanced this paradigm by exploring various architectural variants [5, 40] and extending it to conditional generation tasks [6, 7], particularly point cloud to mesh generation. Most recently, EdgeRunner [33] proposes using half-edge algorithm [29] for efficient tokenization. However, these methods exhibit manifold defects when handling complex topologies and generate excessively long sequences, resulting in a limited number of faces or prohibitively high training costs. In contrast, our Nautilus aims to achieve manifold structure fidelity by a compact, locality-aware mesh tokenization and conditioning mechanism.
@@
Figure 4: The illustration of our Nautilus-style Mesh Tokenization algorithm. (a) The shell-structured traversal of mesh faces. (b) Top: The continuous shell traversal designed to find the center of the next shell. Bottom: The tokenized vertices sequence of the shown example.
3 Nautilus
@@
In this section, we introduce Nautilus, a locality-aware autoencoder for artist-like mesh generation, with the overall pipeline depicted in Figure 3. Initially, we design a novel Nautilus-style algorithm that tokenizes artist-created meshes into sequences (Sec. 3.1), which serves as the training input of our framework. During generation, Nautilus takes either a point cloud or a single-view image as the condition inputs. For point cloud conditioning, we propose a dual-stream conditioner (Sec. 3.2) that provides both global and local geometric guidance to the autoregressive decoder, which generates the tokenized sequence given the input condition (Sec. 3.3). For image conditioning, we integrate a pre-trained image encoder with the decoder to enable image-conditioned generation (Sec. 3.4). Finally, after detokenization, the token sequences predicted by the decoder are translated into mesh faces, constructing our generated artist-like mesh assets.
3.1 Nautilus-style Mesh Tokenization
@@
The initial step of autoregressive mesh generation involves serializing the triangle mesh into tokens [5, 31, 6]. A mesh asset $\mathbf{M}=\{\mathbf{f}^{i}\}_{i=1}^{N}$ consists of $N$ triangle faces $\mathbf{f}^{i}$, each of which is represented by three vertices, $\{\mathbf{v}_{1}^{i},\mathbf{v}_{2}^{i},\mathbf{v}_{3}^{i}\}$. By flattening the discrete 3D coordinates of each vertex, denoted as $\mathbf{v}_{j}^{i}=\{x_{j}^{i},y_{j}^{i},z_{j}^{i}\}$, the mesh asset $\mathbf{M}$ can be serialized into an ordered sequence of coordinates, $S(\mathbf{M})$, with a total length of $9N$:
$S(\mathbf{M})=\{\mathbf{v}_{1}^{1},\mathbf{v}_{2}^{1},\mathbf{v}_{3}^{1},\mathbf{v}_{1}^{2},\mathbf{v}_{2}^{2},\mathbf{v}_{3}^{2},\dots,\mathbf{v}_{1}^{N},\mathbf{v}_{2}^{N},\mathbf{v}_{3}^{N}\}$ (1)
@@
Nautilus Shell Representation. In tokenization, we first sort the mesh faces by their coordinates $z$-$y$-$x$ to facilitate efficient serialization. As depicted in Figure 4(a), we divide the mesh asset into multiple shells for sequential generation. Each shell organizes faces around a central vertex $O$ with an ordered sequence of surrounding vertices $P$, where each face is formed by $O$ and two adjacent $P$ as $\mathbf{f}_{OP_{i}P_{i+1}}$. This organization helps maintain local mesh connectivity by explicitly encoding edge-sharing relationships between adjacent faces. Based on this shell structure, we compactly represent $N$ faces using a sequence of $(N+2)$ vertices, rather than the $3N$ vertices required in Eq. (1):
$S(\mathbf{f}_{OP_{1}P_{2}},\mathbf{f}_{OP_{2}P_{3}},\dots)=\{O,P_{1},P_{2},P_{3},\dots\}$ (2)
@@
With this representation, the shell #1 in Figure 4(a) can be serialized to $\{O_{1},P_{1,1},P_{1,2},\dots,P_{1,10},P_{1,1}\}$, which preserves the adjacency of neighboring vertices in sequence. By explicitly modeling the locally shared vertices and faces, we could add a new face to a shell by extending the sequence with only a single vertex.
@@
Continuous Shell Traversal. After completing the traversal of one shell, we decide the central vertex candidate for the next shell based on two insights: (1) The vertex should be close to the current shell to facilitate local, short-distance dependencies and promote connectivity between shells. (2) The vertex should have a relatively high degree and is shared by more faces, which not only improves the compression ratio but also encourages the formation of well-connected local mesh structures. Therefore, we design the following continuous shell traversal strategy. Specifically, after completing one shell, we list all neighbor vertices of its last traversed vertex and select the neighbor with maximum degree as the center vertex of the next shell  [^1] As depicted in Figure 4(b), the last vertex of the shell #k is located at $P_{k,\text{end}}$, therefore we choose the orange vertex with the highest degree among its neighbors as the next center $O_{k+1}$.. Our strategy enhances the traversal continuity and spatial proximity between shells, promoting locally consistent mesh topology.
@@
Coordinates Compression. Building upon the vertex sharing optimization in our Nautilus representation, we exploit spatial quantization and dimensional reduction techniques to achieve more compact mesh encoding. Specifically, given the 3D spatial resolution $\alpha$, we select a proper integer multiplier $\beta$ and uniquely represent the 3D discrete coordinates $(x,y,z)\in\mathbb{R}^{\alpha}$ in a 2D space $(u,v)$:
$x*\alpha^{2}+y*\alpha+z=u*\beta+v$ (3)
@@
By mapping the $x$-$y$-$z$ coordinates to $u$-$v$ space, we could further compress the sequence length. With a 128 resolution and a multiplier of 2048, we construct a codebook of size 1024 for $u$ and 2048 for $v$. To identify the start of each shell, we also extend the codebook by 1024 to separately encode the $u$ coordinates of center vertices $O$, thereby distinguishing them from surrounding vertices $P$. This design avoids the increased sequence length introduced by the special tokens that previous work [7, 33] used for sub-sequence separation, producing our tokenized sequence as:
$S(\mathbf{M})=\{u^{\text{O}}_{1},v^{\text{O}}_{1},u^{\text{P}}_{1,1},v^{\text{P}}_{1,1},u^{\text{P}}_{1,2},\dots,u^{\text{O}}_{2},v^{\text{O}}_{2},\dots\}$ (4)
@@
In a nutshell, our Nautilus-style tokenization provides a compact serialization to model complex mesh assets with a higher number of faces. Meanwhile, it maintains short-distance local dependencies by preserving the proximity of spatially neighboring vertices in the tokenized sequence, enhancing the fidelity of local structure. Experiments in Sec. 4 show that these advantages lead to significant improvements in scalability and generation quality. A detailed step-by-step demonstration of the tokenization process in Appendix, further illustrates its design and implementation.
3.2 Dual-Stream Point Conditioner